OpenStudy (anonymous):

What is the equation of a line that passes through points (-1, 2) and (5, 2)?

OpenStudy (anonymous):

y=mx+b m=(y2-y1)/(x2-x1) y2=2 y1=2 x2 =5 x1=-1 m=0 y=0x+b y=b y=2